Objects with masses of 250 kg and a 550 kg are separated by 0.490 m. (a) Find the net gravitational force exerted by these objects on a 63.0 kg object placed midway between them. magnitude .000021 Correct: Your answer is correct. N direction Correct: Your answer is correct. (b) At what position (other than infinitely remote ones) can the 63.0 kg object be placed so as to experience a net force of zero? Just need part B
Answer #1

b)

If the distance from the larger mass is x, then the distance from the smaller mass is 0.490 - x. In order for the forces to be equal, then
G * 63 * 250 / (0.490 - x)^2 = G * 63 * 550 / x^2

G and 63 are cancels:

250 * x^2 = 550 * (0.490 - x)^2

by solving this euqaiton

x = 0.29267 m

x = 1.5039 m